Monday, August 28, 2006
- Additional progress in restoring temporary electrical power to 1542 Tulane Ave. was made today.
- The electrical grounds have been made from the 1000 KV and 500 KV transformers to the water pipe and the grid in the vault.
- The electricians finished piping in the 13.8 V transformer to the switchgear. All of the wires have been pulled except for the wires to the vault.
- The electrical grounds are being connected to the 13.8 Amp switchgear.
- The contractor is now tracing the circuits to the PL2 and PL3 panels and the loads are being terminated as each trace is completed.
- The 480 V circuits are landed to the troughs and the starter disconnect.
- The electricians are flagging and painting the 13.8V pipes red.
- In the MEB, the last of the work is underway to make Lecture Room A available for use.
- The data testing subcontractor finished testing the connections at the pods and in the data cabinets over the weekend. The pods and the data cabinet must still be labeled.
- At the Dental School, the removal of flood damaged medical records is now complete. Additional notes on the work include the following:
- The elevator contractors have changed the buffers on Elevators #6 & #7. However, further elevator work is now on hold until additional parts and materials arrive.
- The medical gas system (oxygen & nitrous oxide) is scheduled for testing on Wednesday, August 30. The system is currently pressurized with 70 lbs of pressure & holding.
- The labeling of the new piping and insulation is currently underway. The contractor is beginning with the chilled water piping, which is approximately 80% complete.
- Mold remediation of heavily molded rooms (6413B, 6409, & 7409A) is nearing completion. The rooms did not pass the air tests and must be retested. Retesting is scheduled for later this week.
- The domestic water system successfully passed water testing for sterilization.